What is the explicit formula for this geometric sequence?

f 1, 2, 3, 4, 5,
f(n) 1, 4, 8, 16, 32
1
Expert's answer
2019-09-24T08:00:11-0400

The formula for any geometric sequence is


"b_n=b_1q^{n-1},"

where "b_1" - the first element, "q" - common ratio, "n" - number of an element.

Here we have sequence numbers from 1 to 5, i.e. 5 elements, and 5 values for these sequence elements.

Looking at the equation above, we can notice that according to the condition either there should be 6 elements in f row and as well as we must include 2 into f(n), or it must be 2 instead of 1 in f(n) row. There is no any geometric sequence with the numbers you wrote.

Therefore, assuming case a):


"n\space\space\space\space\space1 \space2\space 3\space 4\space \space5\space \space\space6 \\
f(n)\space1 \space2 \space4 \space8 \space16 \space32\\
b_n=1\cdot2^{n-1}=2^{n-1}."

Assuming case b):


"n\space\space\space\space\space\space1 \space2\space 3\space\space 4\space \space5\space\\
f(n)\space\space2 \space4 \space8 \space16 \space32\\
b_n=2\cdot2^{n-1}=2^n."